OpenStudy (anonymous):

lim of (sin x-1) / (x - pi/2) as x approaches Pi/2

OpenStudy (freckles):

\[\lim_{x \rightarrow \frac{\pi}{2}}\frac{\sin(x)-1}{x-\frac{\pi}{2}}\] recall sin(pi/2)=1\[\lim_{x \rightarrow \frac{\pi}{2}}\frac{\sin(x)-\sin(\frac{\pi}{2})}{x-\frac{\pi}{2}}\] does this look familiar?

OpenStudy (anonymous):

Yes.

OpenStudy (freckles):

\[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=f'(x)|_{x=a}\]

OpenStudy (freckles):

so if f(x)=sin(x) then f'(x)=?

OpenStudy (anonymous):

cos x

OpenStudy (freckles):

and what is cos(x) evaluated at x=pi/2

OpenStudy (anonymous):

right?

OpenStudy (freckles):

yes f'(x)=cos(x) but what is f'(a) where a=pi/2

OpenStudy (anonymous):

0

OpenStudy (freckles):

yep :)

OpenStudy (freckles):

\[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=f'(x)|_{x=a} \\ \lim_{x \rightarrow \frac{\pi}{2}}\frac{\sin(x)-\sin(\frac{\pi}{2})}{x-\frac{\pi}{2}}=\cos(x)|_{x=\frac{\pi}{2}}=\cos(\frac{\pi}{2})=0 \]

OpenStudy (freckles):

and I could have wrote it as f'(a) but for some reason that confuses people

OpenStudy (freckles):

and if you knew l'hosptal you could have done that too

OpenStudy (tkhunny):

f'(a) is a bit of an odd notation. Your version is much more clear. f(a) is a constant. The 1st derivative of f(a) should be zero.

OpenStudy (freckles):

f'(a) doesn't mean f itself is a constant function f'(a) means to find the general form for f'(x) and evaluate x at x=a